A 15,600 kg rocket in space fired two thrusters. One thruster propelled the rocket forward with a force of 77,500 N while the second thruster propelled the rocket to port (left) with a force of 27,200 N. Find its acceleration (angle and magnitude).



1
Expert's answer
2020-12-25T14:07:55-0500


First, let us find the resultant force. Since both forces "\mathbf{F}_1" and "\mathbf{F}_2" (see the figure) are perpendicular to each other, the magnitude of the resultant force "\mathbf{F}" can be found with Pythagorean theorem:

"F = \sqrt{F_1^2 + F_2^2}"

where "F_1 = 77500N" and "F_2 = 27200N" are magnitudes of the forces "\mathbf{F}_1" and "\mathbf{F}_2" respectively. Thus, obtain:


"F = \sqrt{77500^2 +27200^2}\approx 82135N"

The magnitude of the acceleration is (according to the second Newton's law):


"a = \dfrac{F}{m} = \dfrac{82135N}{15600kg} \approx 5.3m/s^2"

where "m = 15600kg" is the mass of the rocket.

The angle of the acceleration with x-axis is equal to the corresponding angle "\theta" (from the right triangle in the figure):


"\theta= \arctan\left( \dfrac{-F_1}{F_2}\right)\\
\theta = 180\degree - \arctan\left( \dfrac{77500}{27200}\right) = 180\degree-70.7\degree = 109.3\degree"

Answer. Magnitude: 5.3 m/s^2, angle: 109.3 degrees.